OpenStudy (anonymous):

What is the simplified sum of square root 5+square root 125

OpenStudy (anonymous):

6sqrt(5) right?

OpenStudy (mertsj):

\[\sqrt{5}+\sqrt{125}=\sqrt{5}+\sqrt{25 \times 5}=\sqrt{5}+\sqrt{25}\sqrt{5}=\sqrt{5}+5\sqrt{5}=6\sqrt{5}\]